# TathyaNyaya and FactLegalLlama: Advancing Factual Judgment Prediction and Explanation in the Indian Legal Context

### v1

@@ 8 Conclusions and Future Work
Another significant limitation was the lack of extensive expert evaluation for the generated explanations. While we used high-quality annotations for the TathyaNyaya dataset, resource-intensive processes required for legal expert reviews made it impractical to evaluate the entire dataset. Instead, evaluations were conducted on a smaller subset, which, while insightful, may not fully represent the model’s performance across diverse legal scenarios.
The model’s performance on scrapped datasets was also not fully evaluated due to configuration constraints, leaving gaps in understanding its generalizability to non-annotated factual data. Furthermore, challenges such as hallucinations in generative outputs and maintaining factual consistency in explanations remain unresolved, which can impact the reliability of the model in real-world legal applications.
Lastly, the dataset used in this study comprises only English-language judgments, which limits its applicability in multilingual contexts, especially in jurisdictions where regional languages dominate legal proceedings. This exclusion highlights the need for more inclusive datasets that reflect the linguistic diversity of legal documents in India and beyond.
@@ A.1 Transformers Training Hyper-parameters
To establish competitive baselines, we fine-tuned transformer models such as InLegalBERT and XLNet_Large on the NyayaFacts dataset. Each model was trained with a batch size of 16 using the AdamW optimizer Kingma and Ba (2014) and a learning rate of 2e-6. We ran the training for three epochs, adopting default hyper-parameter settings from the HuggingFace Transformers library. Experiments were carried out on an NVIDIA A100 40GB GPU, ensuring adequate computational resources for handling extensive legal text. This training protocol allowed the models to capture the nuances of fact-based segments and reliably predict judicial outcomes.
@@ A.4 Training Procedure for Hierarchical BiLSTM-CRF Classifier
### Input: <<<{case_facts}>>>
@@
### Input: <<<{case_facts}>>>
@@
### Input: Paraphrase the following text:<<<{case_facts}>>>

### v2

@@ Acknowledgements
The dataset used in this study comprises only English-language judgments, which limits its applicability in multilingual contexts, especially in jurisdictions where regional languages dominate legal proceedings. This exclusion highlights the need for more inclusive datasets that reflect the linguistic diversity of legal documents in India and beyond.
@@ A.1 Transformers Training Hyper-parameters
To establish competitive baselines, we fine-tuned transformer models such as InLegalBERT and XLNet_Large on the NyayaFacts dataset. Each model was trained with a batch size of 16 using the AdamW optimizer kingma2014adam and a learning rate of 2e-6. We ran the training for three epochs, adopting default hyper-parameter settings from the HuggingFace Transformers library. Experiments were carried out on an NVIDIA A100 40GB GPU, ensuring adequate computational resources for handling extensive legal text. This training protocol allowed the models to capture the nuances of fact-based segments and reliably predict judicial outcomes.
@@ A.4 Training Procedure for Hierarchical BiLSTM-CRF Classifier
### Input: <<{case_facts}>>
@@
### Input: <<{case_facts}>>
@@
### Input: Paraphrase the following text:<<{case_facts}>>